OpenStudy (anonymous):

Solve 2(3 – 2x) = x – 9.

OpenStudy (anonymous):

6 - 4x = x - 9

OpenStudy (anonymous):

6 - 5x = -9

OpenStudy (anonymous):

-5x = -15

OpenStudy (anonymous):

x = 3

OpenStudy (anonymous):

Did I do this problem correctly?

ganeshie8 (ganeshie8):

perfect !

ganeshie8 (ganeshie8):

also you can confirm yourself by putting x=3 in the initial equation and checkiing if it satisfies